7. Harvey Chemicals Pty Ltd manufactures a product called Hartik. Direct materials are added at the beginning of the process and conversion activity occurs uniformly throughout the process. The following data pertains to the month of October.

Using the weighted average method of process costing, calculate the equivalent units of direct materials and conversion costs for the month of October.

A. Direct Material 75 000; Conversion Cost 69,400 B. Direct Material 75 000; Conversion Cost 60,400 C. Direct Material 60 000; Conversion Cost 60,400 D. Direct Material 68 000; Conversion Cost 69,400

2

Questions 8 and 9 relate to the following information.

The Singleton Boot Company has four departments in its factory with two service departments, S1 and S2, and two production departments Assembly and Finishing.

The current costs of each department are:

Service Departments

S1 S2

$308 000 $256 000

Production Departments

Assembly $505 000 Finishing $495 000

The distribution and consumption of services is given in the following table:

Services provided to:

S1

S2

Assembly

Finishing

Services provided by:

S1

-

35%

45%

20%

S2

15%

-

60%

25%

8.Using direct method how much overhead costs will be allocated from S1 to Assembly:

A. $138,600 B. $213,231 C. $115,200 D. $177,231

9.Using step-down method how much overhead costs will be allocated from S2 to Assembly:

A. $153,600 B. $218,280 C. $256,800 D. $180,706

3

10. Nala Corporation uses direct labour-hours in its predetermined overhead rate. At the beginning of the year, the estimated direct labour-hours were 14,400 hours and the total estimated manufacturing overhead was $355,680. At the end of the year, actual direct labour hours for the year were 15,200 hours and the actual manufacturing overhead for the year was $350,680.

Overhead at the end of the year was:

  1. $24,760 underapplied

  2. $24,760 overapplied

  3. $19,760 underapplied

  4. $19,760 overapplied

  1. A collection of costs that are to be allocated to cost objects is referred to as a:

    A. Cost pool
    B. Costallocation C. Directcost
    D. Indirect cost

  2. A special order generally should be accepted if:

    A. Its revenue exceeds allocated fixed costs, regardless of the variable costs associated with the order
    B. Excess capacity exists and the revenue exceeds all variable costs associated with the order.

    C. Excess capacity exists and the revenue exceeds allocated fixed costs D. The revenue exceeds variable costs, regardless of available capacity

  3. For a manufacturing company, the cost of factory insurance is classified as a:

    A. Direct material cost.

B. Direct manufacturing labour cost. C. Manufacturing overhead cost.
D. Period cost.

4

14. Kaiser Ltd. uses the high-low method to estimate the cost function. The information for 2018 is provided below.

Month

Machine hours

Labour costs ($)

January

130

5,124

February

250

7,611

March

180

6,208

April

165

6,300

May

240

9,800

June

80

3,200

July

155

6326

August

190

7,200

September

230

8,138

October

250

8,600

November

260

10,004

December

190

7,400

What is the estimate of Kaiser’s cost function when 120 machine hours are used (to the nearest dollar)?

A. $4,347 B. $4,712 C. $4,744 D. $4,797

15. Crimson Quilters makes decorative comforters, quilted garments, and other products in a small sewing factory. In 2018, the company expects to make 2,000 comforters. With respect to the comforters how would the fabric used to make the comforters be classified?

A. Direct and variable costs B. Direct and fixed costs
C. Indirect and variable costs D. Indirect and fixed costs

5

16. McDonnell Industries estimated manufacturing overhead for the year at $290,000. Manufacturing overhead for the year was underapplied by $12,000. The company applied $235,000 to Work in Process. The amount of actual overhead would have been:

A. $278,000.
B. $247,000.
C. $223,000.
D. None of these

17. The benefits arising from the introduction of activity-based costing are likely to be greater where:

i. Overhead is a large proportion of total cost.
ii. Implementation costs are not high due to the support of advanced IT systems. iii. Batches are of a similar size.

A. i and ii
B. i and iii
C. ii and iii
D. None of the above

18. Lawler Manufacturing Co. has three divisions which are in competition with each other, selling slightly different products. Annual results (including per unit sales)

show:

Sales price per unit Variable cost per unit Total Fixed costs Predicted unit sales

Division A $250
$160 $110,000 10,000

Division B $280
$210 $85,000 5,000

Division C $330

$220 $105,000

3,000

The company believes that if it drops Division B, sales of Division A will increase by
3000 units and sales of Division C will increase by 1000 units. Analysis reveals that $50,000 of fixed costs relating to Division B are avoidable for the company if Division B is closed. The incremental effect on the income of the company if it closes Division B will be:

A. An increase of $30,000 B. An increase of $50,000 C. A decrease of $70,000 D. An increase of $80,000

6

19. When the step-down method is used to allocate support department costs, a common way to select the first support department in the sequence is to choose the support department which:

  1. Obtains the highest yield

  2. Serves no other support department

  3. Serves the greatest number of other support departments

  4. Obtains the highest yield AND serves no other support department

20. Franklin has received a special order for 10,000 units of its product at a special price of $15. The product normally sells for $20 and has the following manufacturing costs:

Per unit Direct materials $6

Direct labour 3 Variable manufacturing overhead 2 Fixed manufacturing overhead 6 Unit cost $17

Franklin is currently operating at full capacity and cannot fill the order without harming normal production and sales. If Franklin accepts the order, what effect will the order have on the company’s short-term profit?

A. $40,000 decrease

  1. $40,000 increase

  2. $50,000 decrease

D. $10,000 increase
Answer #1
8.Direct method of OH cost allocation
All the OH costs of the service depts. Are allocated fully,only among the production depts.
from S1 to Assembly
308000*45%/(45%+20%)=
308000*45%/65%=
213231
ANSWER: B. $ 213231
9. Step-down method Production Depts. Service depts.
Assembly Finishing S1 S2
Costs before allocation 505000 495000 308000 256000
Allocation of S1 costs(being 308000> 256000)(45%,20%& 35%) 138600 61600 -308000 107800
363800
Allocation of total S2 costs(60%,25%) 256800 107000 -363800
Total costs after step-down allocation of Ohs 900400 663600 0 0
So, ANSWER: C. OHs allocated from S2 to Assembly= $ 256800
10.Nala Corporation
Predetermined OH rate= Estimated MOH/Estd. D/L hrs.
ie. $ 355680/14400 D/L hrs.
24.7
per D/L hr.
So, actual D/L hrs. applied will be
Actual D/L hrs.*Predetermined OH rate.hr.
ie. 15200 hrs.* $ 24.7=
375440
Actual MOH= 350680
Actual MOH($ 350680) < MOH applied ($ 375440)
So, MOH over-applied=
375440-350680=
24760
ANSWER: B
11.A collection of costs that are to be allocated to cost objects is referred to as a
A. Cost pool
ie. Costs are accumulated so as to be assigned to different cost objects , such as a product/branch, etc.
12.A special order generally should be accepted if:
B. Excess capacity exists and the revenue exceeds all variable costs associated with the order.
so as to maximise contribution income , fixed costs remaining the same.
13.For a manufacturing company, the cost of factory insurance is classified as a:
C. Manufacturing overhead cost.
It is also a period cost--ie. Which does not vary/change according to units produced, but incurred for the period.
It is a fixed mfg. OH cost, ie. A fixed indirect cost.
14.As per the High-Low method,
Labor costs at maximum activity level 260 10004
Labor costs at minimum activity level 80 3200
Difference 180 6804
Difference in cost/activity diff. 6804/180=
37.8
Now, finding the fixed cost ,
Total cost=(M/c hrs.*37.8)+Fixed costs
at 250 m/c hrs.
ie. 10004=(260*37.8)+FC
solving the above, FC= 176
at 80 m/c hrs.
ie. 3200=(80*37.8)+FC
solving the above, FC is the same = 176
So,now, the cost equation at any level of activity will be
Total cost=(M/c hrs.*37.8)+176
So,estimate of Kaiser’s cost function when 120 machine hours are used (to the nearest dollar) will be:
Total cost=(120*37.8)+176
4712
ANSWER: B. $ 4712
15. A. Direct and variable costs
as fabric used to make the comforters are raw materials for the latter & vary directly with the end product.
16.McDonnell Industries
MOH applied 235000
MOH under-applied 12000
So, actual MOH incurred(235000+12000) 247000
Actual MOH($ 247000) > MOH applied ($ 235000)
So, MOH under-applied
ANSWER: B. $ 247000
17..ANSWER: A. i & ii
i-- as ,thenaccuracy of product costing will be very helpful for pricing of the same
ii--will be cost effective without much outlay.
iii.--may not be the answer--then batch costing will be adopted & ABC implementation is more activity based
18. Lawler Manufacturing Co.
Division A B C Total Division A B C Total
S.P./unit 250 280 330 S.P./unit 250 330
V.C/unit 160 210 220 V.C/unit 160 220
Contn./unit 90 70 110 Contn./unit 90 110
Salesunits 10000 5000 3000 Salesunits 13000 4000
Total contn. 900000 350000 330000 1580000 Total contn. 1170000 0 440000 1610000
FC 110000 85000 105000 300000 FC 110000 35000 105000 250000
Profit 790000 265000 225000 1280000 Profit 1060000 -35000 335000 1360000
Incremental profit 80000
ANSWER: D. An increase of $80,000
19. When the step-down method is used to allocate support department costs, a common way to select the first support department in the sequence is to choose the support department which:
Serves the greatest number of other support departments
(ANSWER)
20.(Franklin) Loss of sale value (20-15)*10000= 50000
ANSWER:
$50,000 decrease